OpenStudy (anonymous):

|1/x|>x Looks simple enough, but really I...nvm >.<

OpenStudy (saifoo.khan):

isnt this the same as yesterday's

OpenStudy (anonymous):

Not really, the x in this one is in inverse

OpenStudy (anonymous):

\[|\frac{1}{x}|>x\] like that ?

OpenStudy (anonymous):

exactly! :)

OpenStudy (anonymous):

this one is probably not so bad, because it is true for all \(x<0\)

OpenStudy (anonymous):

i.e. if x is negative, then it is certainly the case that \(|\frac{1}{x}|>x\) right? so now we only have to check what happens when x is positive, and when x is positive you have \(|\frac{1}{x}|=\frac{1}{x}\)

OpenStudy (anonymous):

so your only real job is to solve \[\frac{1}{x}>x\] for \[\frac{1}{x}-x>0\] \[\frac{1-x^2}{x}>0\]

OpenStudy (anonymous):

we can ignore the denominator, because we know that \(x>0\)

OpenStudy (anonymous):

ahhhh I see what you did there :)

OpenStudy (anonymous):

owh lol @saifoo.khan

OpenStudy (saifoo.khan):

;)

OpenStudy (anonymous):

so only have to solve \[1-x^2>0\] and this is a parabola that opens down, so it is positive between the zeros and negative outside them. therefore it is positive on in interval \((-1,1)\) and so your answer is \((-\infty, 1)\) except not when \(x=0\) so i guess you could write \[(-\infty, 0)\cup (0,1)\]

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

yw